Find the probability that a person tossing 3 coins will get either all heads or all tails for the second time on the fifth toss?



Select one:
a. 0.1069
b. 0.105
c. 0.256
1
Expert's answer
2020-06-16T16:08:02-0400

Number of sample spaces when 3 coins tossed= 8

P(getting all heads or all tails )= 2/8=1/4


To find : probability of occurring all heads or all tails in tosses 1-4 and then again on toss 5


solution: Let x be the number of trials until get either all heads or all tails for the second time , the x is a negative binomial random variable.

Formula:

"b(x;k,p)=\binom{x-1}{k-1})p^{k}q^{x-k},"

x= k , k+1,k+2,.....k=2


P(x=5)=b(5;2,1/4)="\binom{4}{1}" (1/4)2(3/4)3=27/256=0.105

Answer: 0.105